# Humanoid Parkour Learning

### v1

@@ Abstract
Parkour is a grand challenge for legged locomotion, even for quadruped robots, requiring active perception and various maneuvers to overcome multiple challenging obstacles. Existing methods for humanoid locomotion either optimize a trajectory for a single parkour track or train a reinforcement learning policy only to walk with a significant amount of motion references. In this work, we propose a framework for learning an end-to-end vision-based whole-body-control parkour policy for humanoid robots that overcomes multiple parkour skills without any motion prior. Using the parkour policy, the humanoid robot can jump on a 0.42m platform, leap over hurdles, 0.8m gaps, and much more. It can also run at 1.8m/s in the wild and walk robustly on different terrains. We test our policy in indoor and outdoor environments to demonstrate that it can autonomously select parkour skills while following the rotation command of the joystick. We override the arm actions and show that this framework can easily transfer to humanoid mobile manipulation tasks. Videos can be found at https://humanoid4parkour.github.io
@@ 4 Experiment Results
In this section, we first describe the setup for our humanoid robot and the experiment setup in simulation and in the real world. Then, we use three separate experiments to demonstrate the following questions: 1. Is using fractal noise terrain feasible compared to “feet air time” reward or motion reference to make the foot raise 2. Concurrent works use “feet air time” [21, 23] or motion reference to make the walking policy able to walk up and down stairs without vision. Is the vision for challenging terrain necessary? 3. Previous Parkour [3, 4] uses either single-GPU distillation or distill a student policy from scratch. Why does our method require both initializing student policy from the oracle policy and using multi-GPU acceleration?
@@ 4.2 Fractal Noise with No Motion Prior is Effective
Figure 4: Real-world quantitative results. Our parkour policy achieves the best performance in the 4 difficult tasks compared with the blind walking policy and the built-in MPC controller in Unitree H1. We use the user interface in H1’s controller to raise its foot height so that it can overcome some of the obstacles.
We compare the success rate and average moving distance between the oracle policies trained with fractal noise terrain (ours) and with “feet airtime” [21, 23] reward terms in all 10 skills in simulation. We deploy 10 robots for each type of terrain with no fractal noise and put them in the easiest sub-track. Shown in Table 2, the success rate is computed from the number of robots that successfully overcome the terrain with maximum difficulty. The average moving distance is computed from the average distance before the robot falls or stops in a given type of terrain. Shown in Table 2, our method that trains on fractal noise terrain performs better than that using “feet air time” in the reward function, which further illustrates that the “feet air time” term conflicts with the some of locomotion tasks.
@@
Table 2: We compare the success rate and average moving distance in simulation. Each type of terrain is a connected 3 sub-track with linearly increasing difficulty. Each sub-track has a length of 4.8m, which makes the total length of a given terrain 14.4m. (More experiment results for the other obstacles are shown in the appendix.)
@@ Acknowledgments
Thanks to Qiao Sun, Derun Li, Shaoting Zhu, Xin Duan, Hao Fu, Honglei Zhang for the video shoot. We thank Xuxin Cheng, Shaoting Zhu, Xin Duan, and Derun Li for helps on experiments, valuable discussions and supports. This project is supported by Shanghai Qi Zhi Institute and outdoor videos are supported by ShanghaiTech University.
@@ Appendix F Deployment Details
We use the Unitree H1 humanoid robot for our real-world test, which is equipped with an Intel RealSense D435i and an Intel Core i7 12-core NUC onboard. The robot has 19 joints including the arm and waist. We assign different $kp$, $kd$ parameters for each joint as shown in Table 8. We use ROS2 and Cyclone DDS for the communication between the policy and motors. The depth encoder sends its embedding to the recurrent network at 10Hz. The recurrent network receives and sends the target positions of all joints to the motors at 50 Hz. The motors compute the torque based on their built-in PD controller at the rate of 1000Hz.
@@ Appendix H Results Compared with Different Distillation Methods
Compared with previous work using quadruped robot parkour [3, 4], we aim to investigate and answer why multi-GPU acceleration is necessary to distill such a humanoid parkour policy. We use the same oracle policy and distill the student policy under 3 different settings, shown in Table 11 12 13. We distill the student parkour policy using a randomly initialized network. After 24 hours of distillation, the student policy still struggles to walk steadily. We test the distillation from the oracle parkour policy but only one GPU process to collect the labeled trajectory and train the policy. After 24 hours, it only trains on $4.147\times 10^{6}$ transitions compared to $432\times 10^{6}$ transitions in 4-GPUs variants. This one-GPU variation needs more time to perform as well as our student parkour policy. The one-GPU variant is aware of different types of obstacles, but the legs do not raise high enough to overcome these obstacles.

### v2

@@ Abstract
Parkour is a grand challenge for legged locomotion, even for quadruped robots, requiring active perception and various maneuvers to overcome multiple challenging obstacles. Existing methods for humanoid locomotion either optimize a trajectory for a single parkour track or train a reinforcement learning policy only to walk with a significant amount of motion references. In this work, we propose a framework for learning an end-to-end vision-based whole-body-control parkour policy for humanoid robots that overcomes multiple parkour skills without any motion prior. Using the parkour policy, the humanoid robot can jump on a 0.42m platform, leap over hurdles, 0.8m gaps, and much more. It can also run at 1.8m/s in the wild and walk robustly on different terrains. We test our policy in indoor and outdoor environments to demonstrate that it can autonomously select parkour skills while following the rotation command of the joystick. We override the arm actions and show that this framework can easily transfer to humanoid mobile manipulation tasks. Videos can be found at project website. [^1] Correspondence Email: hangzhao@mail.tsinghua.edu.cn
@@ 4 Experiment Results
In this section, we first describe the setup for our humanoid robot and the experiment setup in simulation and in the real world. Then, we use three separate experiments to demonstrate the following questions: 1. Is using fractal noise terrain feasible compared to “feet air time” reward or motion reference to make the foot raise 2. Concurrent works use “feet air time” or motion prior [67] to make the walking policy able to walk up and down stairs without vision. Is the vision for challenging terrain necessary? 3. Previous Parkour [3, 4] uses either single-GPU distillation or distill a student policy from scratch. Why does our method require both initializing student policy from the oracle policy and using multi-GPU acceleration?
@@ 4.2 Fractal Noise with No Motion Prior is Effective
Figure 4: Real-world quantitative results. Our parkour policy achieves the best performance in the 4 difficult tasks compared with the blind walking policy and the built-in MPC controller in Unitree H1. We use the user interface in H1’s controller to raise its foot height so that it can overcome some of the obstacles. We run 10 trails for each testing configuration.
We compare the success rate and average moving distance between the oracle policies trained with fractal noise terrain (ours) and with “feet airtime” [21, 23] reward terms in all 10 skills in simulation. We deploy 10 robots for each type of terrain with no fractal noise and put them in the easiest sub-track. Shown in Table 2, the success rate is computed from the number of robots that successfully overcome the terrain with maximum difficulty. The average moving distance is computed from the average distance before the robot falls or stops in a given type of terrain. Shown in Table 2, our method that trains on fractal noise terrain performs better than that using “feet air time” in the reward function, which further illustrates that the “feet air time” term conflicts with some of the locomotion tasks.
@@
Table 2: We compare the success rate and average moving distance in simulation. Each type of terrain is a connected 3 sub-track with linearly increasing difficulty. Each sub-track has a length of 4.8m, which makes the total length of a given terrain 14.4m. The success rates are averaged across 10 runs for each case. (More experiment results for the other obstacles are shown in the appendix.)
@@ Acknowledgments
Thanks to Qiao Sun, Derun Li, Shaoting Zhu, Xin Duan, Hao Fu, and Honglei Zhang for the video shoot. We thank Xuxin Cheng, Shaoting Zhu, Xin Duan, Derun Li, and Miao Zhuang for their help with experiments, valuable discussions, and support. This project is supported by Shanghai Qi Zhi Institute and outdoor videos are supported by ShanghaiTech University.
@@ Appendix F Deployment Details
We use the Unitree H1 humanoid robot for our real-world test, which is equipped with an Intel RealSense D435i and an Intel Core i7 12-core NUC onboard. The robot has 19 joints including the arm and waist. We assign different $kp$, $kd$ parameters for each joint as shown in Table 8. We use ROS2 and Cyclone DDS for the communication between the policy and motors. The depth encoder sends its embedding to the recurrent network at 10Hz. The recurrent network receives and sends the target positions of all joints to the motors at 50 Hz. The motors compute the torque based on their built-in PD controller at 1000Hz.
@@ Appendix H Results Compared with Different Distillation Methods
Compared with previous work using quadruped robot parkour [3, 4], we aim to investigate and answer why multi-GPU acceleration is necessary to distill such a humanoid parkour policy. We use the same oracle policy and distill the student policy under 3 different settings, shown in Table 11 12 13. We run 20 tails for each setting in simulation. We distill the student parkour policy using a randomly initialized network. After 24 hours of distillation, the student policy still struggles to walk steadily. We test the distillation from the oracle parkour policy but only one GPU process to collect the labeled trajectory and train the policy. After 24 hours, it only trains on $4.147\times 10^{6}$ transitions compared to $432\times 10^{6}$ transitions in 4-GPUs variants. This one-GPU variation needs more time to perform as well as our student parkour policy. The one-GPU variant is aware of different types of obstacles, but the legs do not raise high enough to overcome these obstacles.